# MFC 5.0: An exascale many-physics flow solver

### v3

@@ Abstract
Many problems of interest in engineering, medicine, and the fundamental sciences rely on high-fidelity flow simulation, making performant computational fluid dynamics solvers a mainstay of the open-source software community. A previous work (Bryngelson et al.  Comp. Phys. Comm. (2021)) made MFC 3.0 a published, documented, and open-source solver with numerous physical features, numerical methods, and scalable infrastructure. MFC 5.0 is a marked update to MFC 3.0, including a broad set of well-established and novel physical models and numerical methods and the introduction of GPU and APU (or superchip) acceleration. We exhibit state-of-the-art performance and ideal scaling on the first two exascale supercomputers, OLCF Frontier and LLNL El Capitan. Combined with MFC’s single-accelerator performance, MFC achieves exascale computation in practice. New physical features include the immersed boundary method, $N$-fluid phase change, Euler–Euler and Euler–Lagrange sub-grid bubble models, fluid-structure interaction, hypo- and hyper-elastic materials, chemically reacting flow, two-material surface tension, magnetohydrodynamics (MHD), and more. Numerical techniques now represent the current state-of-the-art, including general relaxation characteristic boundary conditions, WENO variants, Strang splitting for stiff sub-grid flow features, and low Mach number treatments. Weak scaling to tens of thousands of GPUs on OLCF Summit and Frontier and LLNL El Capitan see efficiencies within 5% of ideal to over 90% of their respective system sizes. Strong scaling results for a 16-times increase in device count show parallel efficiencies over 90% on OLCF Frontier. MFC’s software stack has further improved, including continuous integration, which ensures code resilience and correctness through over 300 regression tests; metaprogramming, which reduces code length and maintains performance portability; and code generation for computing chemical reactions.
@@ 1 Introduction
MFC was introduced to the literature and open source community via Bryngelson et al. [1]. This work described the software design and features of MFC 3.0, a CPU-based compressible multi-component flow code (MFC) for the 5-equation diffuse interface methods of Kapila et al. [2] and Allaire et al. [3] and the six equation diffuse interface model of Saurel et al. [4]. MFC 3.0 included other features that are discussed throughout this manuscript. Since MFC 3.0’s release, the MFC community has added a broad set of physical models, numerical methods, GPU offloading capabilities, and performance optimizations. This paper focuses holistically on MFC’s development over the past four years. We present physical models, numerical methods, validity, software robustness and testing, and performance on diverse architecture sets at scale on exascale machines and beyond.
@@ 1.1 The history of MFC
The MFC codebase archaeology starts in the mid-2000s. At this time, MFC was an unnamed solver developed by Eric Johnsen. This code represented multi-component flow with the diffuse interface method, which were described in Johnsen [5] and Johnsen and Colonius [6]. Following this effort, Vedran Coralic performed a major code rewrite and adorned MFC with its name. These features adorned MFC with a multi-dimensional numerical reconstruction treatment and are described in Coralic and Colonius [7]. Jomela Meng added cylindrical coordinates systems and corresponding simulations of high-speed shock–droplet interaction [8]. Kazuki Maeda added an Euler–Lagrange representation of sub-grid bubble dynamics [9] and acoustic wave generation [10]. Kevin Schmidmayer and Spencer Bryngelson added the 5-equation model of Kapila et al. [2] and the 6-equation model of Saurel et al. [4]. These efforts were conducted in the research group of Tim Colonius at the California Institute of Technology. Spencer Bryngelson later led a restructuring of MFC at the Georgia Institute of Technology, which included additional modeling and numerics features. This effort culminated in the open-sourcing of MFC 3.0 in 2020 and is largely described in Bryngelson et al. [1]. Modeling, method, and software scalability and portability capabilities were added in subsequent years, which are discussed in this manuscript.
@@ 1.2 New MFC capabilities
MFC now includes more capable physical models and numerical methods, software resistance through continuous integration and deployment, code coverage testing, and computation acceleration via GPU and APU devices. Physical features include six new models for phase change, non-polytropic sub-grid bubble dynamics, treatment for elastic materials, chemical reactions and combustion, and surface tension. An immersed boundary method that supports complex geometries from analytical level sets or STL files. New numerical methods include generalized characteristic boundary conditions, improved shock capturing with TENO and WENO-Z constructions, Strang splitting for stiff sub-grid dynamics, and treatment for low-Mach number flow. Significant software additions have made MFC a portable and performant solver on CPUs, GPUs, and APUs from various vendors, including Intel, AMD, and NVIDIA. Significant decreases in run time are in part enabled via metaprogramming and static code generation.
@@ 1.3 Use of MFC
MFC has a history of being used for early access programs on flagship supercomputers. At the time of writing, these programs include JSC JUPITER (JUREAP) and the LLNL El Capitan and OLCF Frontier final and early access systems, including LLNL Tioga (AMD MI300A testbed) and OLCF Spock and Crusher (AMD MI100 and MI250X testbeds). Commodity cluster use of MFC is broad, including the ACCESS-CI [11] (formerly XSEDE [12]) clusters (PSC Bridges [13] and Bridges2 [14], SDSC Comet [15] and Expanse [16], Purdue Anvil [17], NCSA Delta [18] and DeltaAI [19], TACC Stampede1–3 [20, 21, 22], among others) university clusters, cloud computer systems (Intel’s AI cloud [23]), and AMD and Cray’s internal systems, among numerous others. MFC is a SPEChpc benchmark candidate, which comprises software that evaluates the performance of supercomputers [24].
@@ 2.1 Multi-component treatment
MFC uses reduced versions of the Baer–Nunziato model [25] and diffuse interface methods to model multiphase flow. Diffuse interface methods model the presence of each species using volume fractions and allow for a small region of smearing between species. Other approaches to multiphase flow modeling require interface tracking, which requires coupled flow solvers and complex mesh management [26]. Alternatively, mesh-free methods like smoothed particle hydrodynamics could be used, but smoothed particle hydrodynamics requires extensive load balancing and careful treatment of boundary conditions [26]. Diffuse interface methods are used to avoid these difficulties.
@@ 2.1.1 Five-equation models
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\bm{u}\otimes\bm{u}+p\bm{I}-\bm{T}^{v})$ $=0,$
@@
of Kapila et al. [2] by assuming that both species share a velocity and pressure, and $i=1,\dots,N_{\mathrm{comp.}}$ where $N_{\mathrm{comp.}}$ is the number of fluid components. With this model, $\rho$, $\bm{u}$, $p$, and $E$ are the mixture density, velocity, pressure, and total energy (per unit mass). The $\alpha_{i}$ and $\rho_{i}$ are the volume fraction and density of component $i$. Viscous terms are introduced via the mixture viscous stress tensor
@@
where $\eta$ is the mixture shear viscosity and the strain rate tensor is
@@
The term $K\nabla\cdot\bm{u}$ ensures thermodynamic consistency and accounts for the expansion and compression of each species in mixture regions. The $K\nabla\cdot\bm{u}$ is necessary to model phenomena like spherical bubble collapse, but it is non-conservative and can lead to numerical instabilities [27]. For some problems, $K\nabla\cdot\bm{u}$ is not necessary, and the model reduces to that of Allaire et al. [3]. The differences between these models are discussed in Schmidmayer et al. [27].
@@ 2.1.2 Six-equation model
The six-equation model of Saurel et al. [4] allows for pressure disequilibrium between phases and can be used to avoid numerical stability issues associated with the five-equation Kapila model while maintaining thermodynamic consistency. For two fluids, the six-equation model is
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\bm{u}+p\bm{I}-\bm{T}^{v})$ $=0.$
@@
Here, $\rho$ and $\bm{u}$ are the mixture density and velocity, $\alpha_{i}$, $\rho_{i}$, $p_{i}$, $e_{i}$, and $\bm{T}_{i}^{v}$ are the volume fraction, density, pressure, internal energy, and species viscous stress tensor of species $i$ and $\mu$ is a pressure relaxation parameter. The interfacial pressure $P_{\mathrm{I}}$ is
@@
where $z_{i}=\rho_{i}c_{i}$ is the phasic acoustic impedance and the phasic speed of sound $c_{i}$ is
@@
The model is closed using the same mixture rules of the five-equation model in section 2.1.1. Following Schmidmayer et al. [27], the six-equation model in the limit of infinite relaxation can represent the same physical processes as the model of Kapila et al. [2], but avoids numerical instability issues that arise from some interface capture schemes.
@@ 2.2 Equation of state
The five- and six-equation models are closed using the stiffened gas equation of state (EOS), which faithfully models liquids and gasses [28]. In its simplest form, the stiffened gas EOS relates the internal energy to the pressure and density of each species $i$ as
@@ 2.3 Cylindrical coordinates
Axisymmetric and cylindrical coordinates are implemented via the strategies of Johnsen [5] and Meng [8]. The singularity at $r\to 0$ is handled by placing the axis at a cell boundary such that the radius is defined at the cell center. Then, the solution in cells 180 degrees from each other to support stencils for reconstruction. Spectral filtering in the azimuthal direction relaxes the strict time-step restrictions induced by the small cells near the axis [29].
@@ 3 Numerical methods
We next describe the method used to solve for the advective and diffusive terms in the multi-component models of MFC. The strategy closely follows that of Bryngelson et al. [1].
@@ 3.1 Finite volume method
The models in MFC are solved using a finite volume method based on the framework of Coralic and Colonius [7]. The model equations take the form
@@
where $\bm{q}$ is the vector of cell-averaged conservative variables, $\bm{s}$ and $\bm{h}(\bm{q})$ are source terms, and $\bm{F}^{x},\ \bm{F}^{y},$ and $\bm{F}^{z}$ are the fluxes in the $x$-, $y$- and $z$-directions. 9 is integrated in space across each cell-center as
$\frac{\partial\bm{q}_{i,j,k}}{\partial t}=\frac{1}{\Delta x_{i}}$ $\left(\bm{F}^{(x)}_{i-1/2,j,k}-\bm{F}^{(x)}_{i+1/2,j,k}\right)+\frac{1}{\Delta y_{j}}\left(\bm{F}^{(y)}_{i,j-1/2,k}-\bm{F}^{(y)}_{i,j+1/2,k}\right)+...$ (10)
$\frac{1}{\Delta z_{k}}\left(\bm{F}^{(z)}_{i,j,k-1/2}-\bm{F}^{(z)}_{i,j,k+1/2}\right)+\bm{s}(q_{i,j,k})-\bm{h}(\bm{q}_{i,j,k})\left(\nabla\cdot\bm{u}\right)_{i,j,k}$
The flux $\bm{F}^{x}(\bm{q}_{i+1/2,j,k})$ is calculated at the center of the face $\bm{A}_{i+1/2,j,k}$ with all other fluxes being calculated similarly. The divergence term is calculated as
[…]
@@ 3.2.1 WENO Reconstructions
The fluxes in finite-volume methods are calculated using the left and right state of each interface by solving the Riemann problem
@@
where superscripts $\mathrm{(L)}$ and $\mathrm{(R)}$ indicate the state variables reconstructed from the left- and right-hand sides of the finite volume interface. The simplest reconstruction assumes that the state variables are piecewise constant in each cell. This assumption gives the reconstruction
$\bm{q}^{\mathrm{(L)}}_{i+1/2,j,k}=\bm{q}_{i,j,k}\quad\text{and}\quad\bm{q}^{\mathrm{(R)}}_{i+1/2,j,k}=\bm{q}_{i+1,j,k}.$
The piecewise constant assumption results in a first-order total variation diminishing (TVD) scheme that obviates oscillations at interfaces but is low-order accurate, resulting in meaningful numerical dissipation. High-order shock- and interface-capturing weighted essentially non-oscillatory (WENO) reconstructions prevent the smearing of continuities and allow for the capture of strong shockwaves and material interfaces.
WENO reconstructions obtain high-order spatial accuracy by considering a convex combination of lower-order reconstructions. A $(2k-1)$-th order WENO reconstruction in the $x$-direction uses $k$ staggered candidate polynomials to reconstruct the state variable $f_{i+1/2,j,k}$ with $f^{\mathrm{(r)}}_{i+1/2,j,k}$ for $r=0,1,\dots,k-1$. The weighted sum of candidate polynomials,
$f_{i+1/2,j,k}=\sum_{r=0}^{k-1}\omega_{r}f^{\mathrm{(r)}}_{i+1/2,j,k},$
gives the reconstructed state variable $f_{i+1/2,j,k}$. The nonlinear weights $\omega_{r}$ are calculated via the ideal weights $d_{r}$ and smoothness indicators $\beta_{r}$. The reconstruction of $f_{i,j+1/2,k}$ and $f_{i,j,k+1}$ in the $y$- and $z$-directions are calculated using the same procedure. Fifth-order accuracy is maintained via mapped weights, following Henrick et al. [30]. WENO reconstructions can cause spurious oscillations near material interfaces because they are not TVD for such advection equations. Spurious oscillations are avoided by reconstructing the primitive variables, $\{\alpha\rho_{i},\bm{u},p,\alpha_{i}\}$, intead of the conservatives ones $\{\alpha\rho_{i},\rho\bm{u},E,\alpha_{i}\}$ where $i$ is enumerated over the number of components.

### v4

@@ Abstract
Many problems of interest in engineering, medicine, and the fundamental sciences rely on high-fidelity flow simulation, making performant computational fluid dynamics solvers a mainstay of the open-source software community. Previous work MFC 3.0 was made a published, documented, and open-source solver via Bryngelson et al. Comp. Phys. Comm. (2021) with numerous physical features, numerical methods, and scalable infrastructure. MFC 5.0 is a significant update to MFC 3.0, featuring a broad set of well-established and novel physical models and numerical methods, as well as the introduction of GPU and APU (or superchip) acceleration. We exhibit state-of-the-art performance and ideal scaling on the first two exascale supercomputers, OLCF Frontier and LLNL El Capitan. Combined with MFC’s single-accelerator performance, MFC achieves exascale computation in practice, and achieved the largest-to-date public CFD simulation at 200 trillion grid points as a 2025 ACM Gordon Bell Prize finalist. New physical features include the immersed boundary method, $N$-fluid phase change, Euler–Euler and Euler–Lagrange sub-grid bubble models, fluid-structure interaction, hypo- and hyper-elastic materials, chemically reacting flow, two-material surface tension, magnetohydrodynamics (MHD), and more. Numerical techniques now represent the current state-of-the-art, including general relaxation characteristic boundary conditions, WENO variants, Strang splitting for stiff sub-grid flow features, and low Mach number treatments. Weak scaling to tens of thousands of GPUs on OLCF Summit and Frontier and LLNL El Capitan achieves efficiencies within 5% of ideal to over 90% of their respective system sizes. Strong scaling results for a 16-times increase in device count show parallel efficiencies over 90% on OLCF Frontier. MFC’s software stack has undergone further improvements, including continuous integration, which ensures code resilience and correctness through over 300 regression tests; metaprogramming, which reduces code length while maintaining performance portability; and code generation for computing chemical reactions.
@@
Program title: MFC 5.0
Developer’s repository link: https://github.com/MFlowCode/MFC
Licensing provisions: MIT license (MIT)
Programming language: Fortran08 and Python
Journal Reference of previous version: doi: 10.1016/j.cpc.2020.107396
Does the new version supersede the previous version?: Yes
Reasons for the new version: An abundance of new features and capabilities that are challenging to convey outside of an academic article.
Nature of problem: Simulation of compressible flows requires careful physical model selection and treatment of spatial derivatives to keep solutions physically consistent and free of numerical artifacts. The associated methods should be high-order accurate to reduce computational cost.
Solution method: The present software uses multiple physical models and numerical schemes for the treatment of multi-phase, multi-component, chemically reacting, and magnetohydrodynamic flows. Additional physical effects and models are included. The numerical method is based on the finite volume method.
@@ 1 Introduction
MFC was introduced to the literature and open source community via bryngelson21, which described the software design and features of MFC 3.0, a CPU-based compressible multi-component flow code (MFC) for the 5-equation diffuse interface methods of kapila2001 and allaire2002 and the six equation diffuse interface model of saurel2009. MFC 3.0 included other features that are discussed throughout this manuscript. Since MFC 3.0’s release, the MFC community has added a broad set of physical models, numerical methods, GPU offloading capabilities, and performance optimizations. This paper provides a holistic overview of MFC’s development over the past four years. We present physical models, numerical methods, validity, software robustness and testing, and performance on diverse architecture sets at scale on exascale machines and beyond.
@@ 1.1 The history of MFC
The MFC codebase archaeology dates back to the mid-2000s. At this time, MFC was an unnamed solver developed by Eric Johnsen. This code represented a multi-component flow using the diffuse interface method, as described in johnson2008 and johnsen2006implementation. Following this effort, Vedran Coralic performed a major code rewrite and adorned MFC with its name. These features adorned MFC with a multi-dimensional numerical reconstruction treatment and are described in Coralic2014. Jomela Meng added cylindrical coordinate systems and corresponding simulations of high-speed shock–droplet interaction [meng2016]. Kazuki Maeda added an Euler–Lagrange representation of sub-grid bubble dynamics [maeda2018eulerian] and acoustic wave generation [Maeda2017]. Kevin Schmidmayer and Spencer Bryngelson added the 5-equation model of kapila2001 and the 6-equation model of saurel2009. These efforts were conducted in the research group of Tim Colonius at the California Institute of Technology. Spencer Bryngelson later led a restructuring of MFC at the Georgia Institute of Technology, which included additional modeling and numerical features. This effort culminated in the open-sourcing of MFC 3.0 in 2020 and is largely described in bryngelson21. Modeling, method, and software scalability and portability capabilities were added in subsequent years, which are discussed in this manuscript.
@@ 1.2 New MFC capabilities
To make high-fidelity CFD a practical instrument for today’s many-physics problems, a solver must span liquids, gases, and solids (at least) and accommodate complex geometries and broad separations of time and length scales. At the same time, this should be accomplished efficiently on modern accelerators, which serve as the backbone of current supercomputing centers. MFC 5.0 is architected around that premise. The new capabilities unify established and novel models. Each of these capabilities is accompanied by state-of-the-art numerics and GPU/APU offloading, allowing a single code to transition from low-Mach, geometry-rich flows to shock- and detonation-dominated regimes at exascale. In this sense, the new capabilities were crafted to establish an infrastructure for credible, end-to-end simulations across engineering, medicine, and the sciences.
MFC now includes more capable physical models and numerical methods, software resistance through continuous integration and deployment, code coverage testing, and computation acceleration via GPU and APU devices. With these, MFC conducted the largest publicly known CFD simulation at 200T grid points, or 1 quadrillion degrees of freedom [wilfongGB25], and was an ACM 2025 Gordon Bell Prize Finalist. Physical features include six new models for phase change, non-polytropic sub-grid bubble dynamics, treatment for elastic materials, chemical reactions and combustion, and surface tension. An immersed boundary method that supports complex geometries from analytical level sets or STL files. New numerical methods include generalized characteristic boundary conditions, improved shock capturing with TENO and WENO-Z constructions, Strang splitting for stiff sub-grid dynamics, and treatment for low-Mach number flow. Significant software additions have made MFC a portable and performant solver on CPUs, GPUs, and APUs from various vendors, including Intel, AMD, and NVIDIA. Decreases in runtime are, in part, enabled via metaprogramming and static code generation.
@@ 1.3 Use of MFC
MFC has a history of being used for early access programs on flagship supercomputers. At the time of writing, these programs include JSC JUPITER (JUREAP) and the LLNL El Capitan and OLCF Frontier final and early access systems, including LLNL Tioga (AMD MI300A testbed) and OLCF Spock and Crusher (AMD MI100 and MI250X testbeds). Commodity cluster use of MFC is broad, including the ACCESS-CI [access-ci] (formerly XSEDE [Towns2014-po]) clusters (PSC Bridges [PSCBridges] and Bridges2 [Bridges2], SDSC Comet [Comet2015] and Expanse [Expanse2021], Purdue Anvil [Anvil2022], NCSA Delta [Delta] and DeltaAI [DeltaAI], TACC Stampede1–3 [Stampede1, Stampede2, Stampede3], among others) university clusters, cloud computer systems (Intel’s AI cloud [IntelAICloud]), and AMD and Cray’s internal systems, among numerous others. MFC is a SPEChpc benchmark candidate, which comprises software that evaluates the performance of supercomputers [spechpc]; it is also used as part of the OLCF Test Harness for the performance and robustness of OLCF Frontier [wilfong-hpctests].
@@ 2.1 Multi-component treatment
MFC uses reduced versions of the Baer–Nunziato model [ANDRIANOV2004] and diffuse interface methods to model multiphase flow. The core of the methodology, as described in this subsection, remains unchanged from MFC 3.0 [bryngelson21] and readers are directed there for a more fundamental understanding and detailed algorithmic strategy. Changes to the fundamental diffuse interface methodology are limited, except for performance improvements, code testing strategies, and additional physical features that are now accommodated. These additional features are discussed in section 4.
@@ 2.1.1 Five-equation models
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\otimes\bm{u}+p\bm{I}-\bm{T}^{v})$ $=0,$
@@
of kapila2001 by assuming that both species share a velocity and pressure, and $i=1,\dots,N_{\mathrm{comp.}}$ where $N_{\mathrm{comp.}}$ is the number of fluid components. With this model, $\rho$, $\bm{u}$, $p$, and $E$ are the mixture density, velocity, pressure, and total energy (per unit mass). The $\alpha_{i}$ and $\rho_{i}$ are the volume fraction and density of component $i$. Viscous terms are introduced via the mixture viscous stress tensor
@@
where $\eta$ is the mixture shear viscosity, and the strain rate tensor is
@@
The term $K\nabla\cdot\bm{u}$ ensures thermodynamic consistency and accounts for the expansion and compression of each species in mixture regions. The $K\nabla\cdot\bm{u}$ is necessary to model phenomena like spherical bubble collapse, but it is non-conservative and can lead to numerical instabilities [schmidmayer2020]. For some problems, $K\nabla\cdot\bm{u}$ is not necessary, and the model reduces to that of allaire2002. The differences between these models are discussed in schmidmayer2020.
@@ 2.1.2 Six-equation model
The six-equation model of saurel2009 allows for pressure disequilibrium between phases and can be used to avoid numerical stability issues associated with the five-equation Kapila model while maintaining thermodynamic consistency. For two fluids, the six-equation model is
@@
$\frac{\partial\rho\bm{u}}{\partial t}+\nabla\cdot(\rho\bm{u}\otimes\bm{u}+p\bm{I}-\bm{T}^{v})$ $=0.$
@@
Here, $\rho$ and $\bm{u}$ are the mixture density and velocity, $\alpha_{i}$, $\rho_{i}$, $p_{i}$, $e_{i}$, and $\bm{T}_{i}^{v}$ are the volume fraction, density, pressure, internal energy, and species viscous stress tensor of species $i$, and $\mu$ is a pressure relaxation parameter. The interfacial pressure $P_{\mathrm{I}}$ is
@@
where $z_{i}=\rho_{i}c_{i}$ is the phasic acoustic impedance, and the phasic speed of sound $c_{i}$ is
@@
The model is closed using the same mixture rules of the five-equation model in section 2.1.1. Following schmidmayer2020, the six-equation model in the limit of infinite relaxation can represent the same physical processes as the model of kapila2001, but avoids numerical instability issues that arise from some interface capture schemes.
@@ 2.2 Equation of state
The five- and six-equation models are closed using the stiffened gas equation of state (EOS), which accurately models liquids and gases [menikoff1989]. In its simplest form, the stiffened gas EOS relates the internal energy to the pressure and density of each species $i$ as
@@ 2.3 Cylindrical coordinates
Axisymmetric and cylindrical coordinates are implemented via the strategies of johnson2008 and meng2016. The singularity at $r\to 0$ is handled by placing the axis at a cell boundary such that the radius is defined at the cell center. Then, the solution in cells 180 degrees apart from each other supports stencils for reconstruction. Spectral filtering in the azimuthal direction relaxes the strict time-step restrictions induced by the small cells near the axis [mohseni2000].
@@ 3 Numerical methods
We next describe the method used to solve for the advective and diffusive terms in the multi-component models of MFC. The strategy employed here aligns with that of bryngelson21, and the reader is directed to that work for further details.
@@ 3.1 Finite volume method
The models in MFC are solved using a finite volume method based on the framework of Coralic2014. The model equations take the form
@@
where $\bm{q}$ is the vector of cell-averaged conservative variables, $\bm{s}$ and $\bm{h}(\bm{q})$ are source terms, and $\bm{F}^{x},\ \bm{F}^{y},$ and $\bm{F}^{z}$ are the fluxes in the $x$-, $y$- and $z$-directions. 9 is integrated in space as
$\frac{\partial\bm{q}_{i,j,k}}{\partial t}=\frac{1}{\Delta x_{i}}$ $\left(\bm{F}^{(x)}_{i-1/2,j,k}-\bm{F}^{(x)}_{i+1/2,j,k}\right)+\frac{1}{\Delta y_{j}}\left(\bm{F}^{(y)}_{i,j-1/2,k}-\bm{F}^{(y)}_{i,j+1/2,k}\right)$ (10)
$+\frac{1}{\Delta z_{k}}\left(\bm{F}^{(z)}_{i,j,k-1/2}-\bm{F}^{(z)}_{i,j,k+1/2}\right)+\bm{s}(q_{i,j,k})-\bm{h}(\bm{q}_{i,j,k})\left(\nabla\cdot\bm{u}\right)_{i,j,k}.$
The flux $\bm{F}^{x}(\bm{q}_{i+1/2,j,k})$ is calculated at the center of the finite volume face, and the other coordinate directions follow in the same fashion.
@@ 3.2.1 WENO Reconstructions
The fluxes in finite-volume methods are calculated using the left and right states of each interface by solving the Riemann problem
@@
where superscripts $\mathrm{(L)}$ and $\mathrm{(R)}$ indicate the state variables reconstructed from the left- and right-hand sides of the finite volume interface. WENO reconstructions obtain high-order spatial accuracy by considering a convex combination of lower-order reconstructions. A (2k1)(2k-1)-th order WENO reconstruction in the $x$-direction uses $k$ staggered candidate polynomials to reconstruct the state variable $f_{i+1/2,j,k}$ with $f^{\mathrm{(r)}}_{i+1/2,j,k}$ for $r=0,1,\dots,k-1$. The weighted sum of a set of crafted candidate polynomials gives the reconstructed state variables. Fifth-order accuracy is maintained through mapped weights, as described by hendrick2005.